OpenStudy (anonymous):

What is the lateral surface area of a cone with radius 19 cm and slant height 11 cm?

OpenStudy (anonymous):

\[ \pi \times 19 \times 11 \]

OpenStudy (anonymous):

thank you, i thought that was how you figured it out but i just wanted to make sure.

OpenStudy (anonymous):

Glad to help.